A car of mass 1000kg accelerates from rest for 90s in a straight line to reach a sped of 59km/h. After that, it accelerates uniformly at 6m/s^2 for a further 30s.

a) Calculate the speed of the car at the end of the 30s period.
b) Calculate the total distance covered by the car.
1
Expert's answer
2020-08-10T19:54:39-0400

(a)

"v_f=v_i+at"

"v_f=59\times \frac{1000\:\rm m}{3600\:\rm s}+6\:\rm m/s^2\times 30\: s=196.4\:\rm m/s"

(b)

"d=d_1+d_2\\
=\frac{0+59\times \frac{1000\:\rm m}{3600\:\rm s}}{2}\times 90+\frac{59\times \frac{1000\:\rm m}{3600\:\rm s}+196.4\:\rm m/s}{2}\times 30\\
=3929\:\rm m"